OpenStudy (anonymous):

1. what is (3x^5-15x^4+4x^3+11x^2-9x+2) divided by (x^2-5x+2) A.3x^3- 2x+1 B. 3x^3-2x^2+7 C. 3x^3-2x^2+7x+26 D. 3x^3-30x^2+160x-849

OpenStudy (mathmath333):

\[\Huge \begin{array}l\color{red}{\text{w}}\color{orange}{\text{e}}\color{#E6E600}{\text{l}}\color{green}{\text{c}}\color{blue}{\text{o}}\color{purple}{\text{m}}\color{purple}{\text{e}}\color{red}{\text{ }}\color{orange}{\text{t}}\color{#E6E600}{\text{o}}\color{green}{\text{ }}\color{blue}{\text{o}}\color{purple}{\text{p}}\color{purple}{\text{e}}\color{red}{\text{n}}\color{orange}{\text{s}}\color{#E6E600}{\text{t}}\color{green}{\text{u}}\color{blue}{\text{d}}\color{purple}{\text{y}}\color{purple}{\text{ }}\color{red}{\text{!}}\color{orange}{\text{}}\end{array} \]

OpenStudy (anonymous):

Thanks! Do you know the answer? Or can you help me find it?

OpenStudy (mathmath333):

yes i will help just wait few minutes

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

I will fan/medal or whatever you want. I don't really know how it works but I will if you answer

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{ \begin{array}{r}\underline{3x^3-2x+1}\\x^2-5x+2|~~ {\overline{3x^5-15x^4+4x^3+11x^2-9x+2}}\\~\\ -\underline{\left(3x^5-15x^4+6x^3+0x^2-0x+0\right)}\hspace{1.5em}\\ -2x^3+11x^-9x+2\hspace{.33em}\\ -\underline{(-2x^3+10x^-4x+0)}\\ x^2-5x+2\hspace{.33em}\\ -\underline{(x^2-5x+2)}\\ 0\hspace{.33em}\\ \\ \end{array} }\end{align}\)

OpenStudy (anonymous):

Did you divide 3x^5-15x^4+4x^3+11x^2-9x+2 by x^2-5x+2?

OpenStudy (anonymous):

I am having trouble reading what you wrote down, sorry, haha

OpenStudy (mathmath333):

yes of course

OpenStudy (anonymous):

Oh, LONG DIVISION, sorry. Yes I see what you did now, I don't know what I was thinking. Thank you!

OpenStudy (anonymous):

Can you help me with a few others then? I have quite a few. If not that's okay.

OpenStudy (mathmath333):

yes bring it on

OpenStudy (anonymous):

I have nine, but I will give you one at a time. Which expression represents the result of this subtraction? 3x-1/x+2 - x-2/x-1

OpenStudy (anonymous):

Let me type the answers really fast

OpenStudy (anonymous):

A.) 2x+1/ 3 B.) 2x+1 / x^2+x-2 C.) 3x^2-4x+5 / 3 D.) 2x^2-4x+5 / x^2+x-2

OpenStudy (anonymous):

The 3x-1 / x+2 is subtracted by the x-2 / x-1

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{ \dfrac{3x-1}{x+2 }- \dfrac{x-2}{x-1} \hspace{.33em}\\~\\ \normalsize \text{ find the LCD} \hspace{.33em}\\~\\ =\dfrac{(3x-1)(x-1)}{(x+2)(x-1) }- \dfrac{(x-2)(x+2)}{(x+2)(x-1)} \hspace{.33em}\\~\\ =\dfrac{(3x-1)(x-1)-(x-2)(x+2)}{(x+2)(x-1) } \hspace{.33em}\\~\\ }\end{align}\) see and try to solve it